I'm so confused on this problem.... Please help!!
Solve the following quadratic equation by completing the square. Simplify the solutions and write them in
a + bi form
5x^2 + x = −5

Answer by stanbon(75887)


5x^2 + x = −5
----
5(x^2 + (1/5)x + ?) = -5 + 5*?
----
5(x^2 + (1/5)x + (1/10)^2) = -5 + 5(1/10)^2
----
5(x+(1/10))^2 = (-500/100) + (5/100)
-------------------
5(x+(1/10))^2 = -495/100
--
(x+(1/10))^2 = -99/100
----
x + (1/10) = +-(9/10)sqrt(10)i
-----
x = (-1/10) +-(9/10)sqrt(10)i
==================================
Cheers,
Stan H.
=================